OpenStudy (anonymous):

PLEASE HELP!!!!!!!! give the exact value of sin(cos^−1(e^(ln(3√/2)

OpenStudy (anonymous):

eq editor please

OpenStudy (anonymous):

\[\sin (\cos^{-1} (e ^{\ln(\sqrt{3}/2)})\]

OpenStudy (jdoe0001):

\(\bf \huge sin \left(cos^{-1} \left(e^{ln(\frac{\sqrt{3}}{2})}\right)\right)\)

OpenStudy (anonymous):

hint:\[e ^{\ln \left( \frac{ \sqrt{3} }{ 2 } \right)}=\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (jdoe0001):

ahh

OpenStudy (anonymous):

@pgpilot326 thanks i got that part i just dont know how to do the sin cos inverse thing....

OpenStudy (anonymous):

what angle gives yo a cos of \[\frac{ \sqrt{3} }{ 2 }?\]

OpenStudy (anonymous):

@pgpilot326 pi/6 i think

OpenStudy (anonymous):

Remember that the inverse trig functions will give you angles and take Lengths for input. Trig functions take angles for input and give you lengths! You can use this to find the angle from the unit circle as this is one that exists on it. :D Did I give you a link to my trig sheet last time ses11?

OpenStudy (anonymous):

@agentc0re yes thank you! it is very helpful!! so my final answer will be an angle right?

OpenStudy (anonymous):

yep for 0<angle<pi. what are the restrictions on the range of cos ^-1?

OpenStudy (jdoe0001):

you're being asked for the sin(), thus the sine value

OpenStudy (anonymous):

so find sin of pi/6 and your done.

OpenStudy (anonymous):

oh ok thank you all!!!

OpenStudy (anonymous):

@pgpilot326 wait sorry... one more question how does the cosine change to the inverse? (does it change the answer pi/6?)

OpenStudy (anonymous):

pi/6 is not the answer to the whole question, just this part: \[\cos^{-1} \left( \frac{ \sqrt{3} }{ 2 } \right)=\frac{ \pi }{ 6 }\] you still need to find \[\sin \left( \frac{ \pi }{ 6 } \right)\] that's the answer

OpenStudy (anonymous):

ok thanks i was just wondering if the restrictions changed the pi/6 for the first half of the question but i figured it out and the final answer should be 1/2 I think

OpenStudy (anonymous):

i'm pretty sure that \[0 \le \cos^{-1} x \le \] that covers all of the values of cos without repeating.

OpenStudy (anonymous):

what happened to the pi?